OpenStudy (sabrina1):

Find an equation for the linear function which has f( 100 ) = 1000 and f( 350 ) = 4000 .

OpenStudy (anonymous):

f(x) = a+bx

OpenStudy (anonymous):

a+100b =1000 -- (1)

OpenStudy (anonymous):

a+350b=4000 --(2)

OpenStudy (anonymous):

(2)-(1) ---> 250b = 3000 --> b= 12

OpenStudy (anonymous):

so a= 1000 - 100(12) = 1000-1200 = -200

OpenStudy (anonymous):

f(x) = -200 +12x

OpenStudy (sabrina1):

thanks youve been a great help:)